OpenStudy (anonymous):

Which literal equations are equivalent to F = ma? Choose all answers that are correct. A. B. C. D.

OpenStudy (texaschic101):

F = ma (1) divide both sides by " a " and you get : F/a = m (2) divide both sides by " m " and you get : F/m = a